OpenStudy (anonymous):

For each subset of C1 given below, either prove that it is a subspace or show why it is not a subspace. (a) { f(x) belongs to C1 | f'(pi) = 0 } (b) { f(x) belongs to C1 | f'(0) = pi }

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

you need to check vector space axioms. in particular, the second one is not closed under vector addition

OpenStudy (anonymous):

I want to proof each one and solution steps please

OpenStudy (anonymous):

oh wait it says the derivative right? ok i still think it is not closed under addition

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok to prove something is not a vector space, you need only to show that one example violates one axiom

OpenStudy (anonymous):

ok I need steps for the proof Can you give me steps? , please

OpenStudy (anonymous):

so for example if \[f(x)=x^2+\pi x, g(x)=2x^2+\pi x\] then \[f'(x)=2x+\pi, g'(x)=4x+\pi\] and so \[f'(0)=g'(0)=\pi\] but \[(f+g)'(0)=2\pi\neq \pi\]

OpenStudy (anonymous):

if you choose as a function y=constant,it's lie in C1 but not pass throught 0.

OpenStudy (anonymous):

satelite73 --> is it part b?

OpenStudy (anonymous):

yes that is my example for part b, as to why that is not a vector space, since it is not closed under addition

OpenStudy (anonymous):

"0" must be in subsapace

OpenStudy (anonymous):

ok I see Thanks, do you have an example for the first one?

OpenStudy (anonymous):

what raphaelF said is quicker

OpenStudy (anonymous):

i believe the first one is. you can check the axioms zero is certainly in it, and it is also closed. most of the other properties are inherited from the definintion of addition of funcitons and multiplying a function by a constant

OpenStudy (anonymous):

Can you explain a little bit more

OpenStudy (anonymous):

??

OpenStudy (anonymous):

do you know the axioms you must check?

OpenStudy (anonymous):

addition - multiply - zeros

OpenStudy (anonymous):

I think there are 10

OpenStudy (anonymous):

here is a succinct list http://tutorial.math.lamar.edu/Classes/LinAlg/VectorSpaces.aspx

OpenStudy (anonymous):

for example \(f,g\) have the property that \(f'(\pi)=g'(\pi)=0\) then since \((f+g)'=f'+g'\) if follows that \((f+g)'(\pi)=0\) this shows it is closed

OpenStudy (anonymous):

i think the first is not too: \[f(x)=1\rightarrow f'(x)=0\] ,but f(x) not pass throught zero

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

So both of them not a subspaces riight?

OpenStudy (anonymous):

i think i will respectfully disagree. or at least disagree that because a constant function is in the set that that doesn't make is a vector space. that would disqualify most spaces of functions

OpenStudy (anonymous):

I don't understand that

OpenStudy (anonymous):

the function does not need to take on the values 0 to be in a vector space. it is rather that the space of functions must include the zero function, which this one does

OpenStudy (anonymous):

check the axioms one by one to see if your first set is a vector space. i think that you will find that it is that is the point of this exercise, to learn how to check the axioms

OpenStudy (anonymous):

What do you mean "it's closed"?

OpenStudy (anonymous):

if \(f\in V\) and \(g\in V\) then \(f+g\in V\)

OpenStudy (anonymous):

So the first one is subspace and the second one is not?

OpenStudy (anonymous):

that is what i think, yes. but you should still check the axioms

OpenStudy (anonymous):

some you don't really have much to check, like the associative law

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

lets check one, and lets call our vector space V lets check that for a scalar c, cf is in V since \((cf)'=cf'\) it follows that if\(f\in V\) i.e. if \(f'(\pi)=0\) then likewise \((cf)'(\pi)=0\) and so V is closed under scalar multiplication

OpenStudy (anonymous):

yes right